The Brenmar Sales Company had a gross profit margin​ (gross profitsdivided by​sales) of 25 percent and sales of $ 9.7 million last year.  76 percent of the​ firm's sales are on​ credit, and the remainder are cash sales. ​ Brenmar's current assets equal $ 1.8 ​million, its current liabilities equal $ 295 comma 200​, and it has $ 103 comma 800 in cash plus marketable securities.

a. If​ Brenmar's accounts receivable equal $ 562 comma 500​, what is its average collection​ period?

b. If Brenmar reduces its average collection period to 25 ​days, what will be its new level of accounts​ receivable?

c. ​Brenmar's inventory turnover ratio is 8.8 times. What is the level of​ Brenmar's inventories?
Answer #1

1.
=365*Accounts Receivable/Net Credit Sales
=365*562500/(9.7*10^6*76%)
=27.85031199

2.
=25/365*9.7*10^6*76%
=504931.5068

3.
=Sales*(1-gross profit margin)/Inventory turnover
=9.7*10^6*(1-25%)/8.8
=826704.5455